OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 3x - y = -3 and passes through the point (4, -2). y = -one thirdx - 6 y = -one thirdx - 14 y = 3x - 6 y = 3x - 14

OpenStudy (anonymous):

@Preetha @perl @dan815

OpenStudy (welshfella):

first rearrange the equation to the for y = mx + b where m = slope

OpenStudy (welshfella):

3x - y = -3 -y = -3x -3 multiply thru by -1:- y = 3x + 3

OpenStudy (anonymous):

but 3 is not one of the intercepts

OpenStudy (welshfella):

the slope of the line is 3 so the slope of a line parallel to it will have the same slope 3

OpenStudy (welshfella):

- i am only dealing with the slope at the moment

OpenStudy (welshfella):

to find our equation we can use the standard form y-y1 = m(x - x1) where m = slope = 3 in our case and (x1,y1) is a point on the line x1 = 4 and y1 = -2

OpenStudy (welshfella):

can you continue - plug those values into the standard form

OpenStudy (anonymous):

thanks

OpenStudy (welshfella):

yw the answer is obviously either c or d